OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) + 6.

thomaster (thomaster):

first get rid of the parenthesis

OpenStudy (anonymous):

Okay so then it will be 3x+3= -2x+1+6?

thomaster (thomaster):

Almost -2(x-1)+6 will become -2x+2+6

thomaster (thomaster):

So you'd have 3x+3= -2x+8

OpenStudy (anonymous):

Then subtract 3 and 8?

thomaster (thomaster):

you subtract 3 on both sides to get the numbers to the right

OpenStudy (anonymous):

so it will be 3x = -2+5?

OpenStudy (anonymous):

-2x

thomaster (thomaster):

3x=-2x+5 Now you need to get the -2x to the left To do that you add 2x on both sides

OpenStudy (anonymous):

5x=5 then you divide it and it equals 1?

thomaster (thomaster):

yes you divide both sides by 5 so the result would be x=1 Well done :)

OpenStudy (anonymous):

Thank you (:

thomaster (thomaster):

You're welcome